OpenStudy (ammarah):

Solve the system using any algebraic method.. x-2y+4z= -19 2x+y-3z+14 3x+y + 2z=5

OpenStudy (ammarah):

@campbell_st

OpenStudy (campbell_st):

looking at it the easiest to eliminate is y double the 2nd equation and add it to the 1st equation double the 3rd equation and add it to the 1st equation. now you will have 2 equations in 2 unknowns.... you can use elimination or substitution to solve for x and z... when you get them substitute into any of the original equations to find y hope this helps

OpenStudy (ammarah):

can u pleasse show me how to do this prob....i know the directions but i still dont get it until i see how its done..

OpenStudy (wolf1728):

I'm guessing that 2x+y-3z+14 means 2x+y-3z = +14

OpenStudy (ammarah):

yeah sorry

OpenStudy (campbell_st):

ok... double equation 2 and add to equation 1 to eliminate y x - 2y + 4z = -19 + 4x + 2y -6z = 28 ------------------ 5x -2z = 9 double equation 3 and add equation 1 x - 2y + 4z = -19 + 6x + 2y + 4z = 10 -------------------- 7x + 8z = -9 so now you need to solve 5x -2z = 9 7x + 8z = -9 so multiply the 1st equation by 4 and add to the 2nd equation, this will eliminate z 20x - 8z = 36 + 7x + 8z = -9 ----------------- 27x = 27 solve for x hope this make sense

OpenStudy (ammarah):

i alsoneed to find y and z

OpenStudy (campbell_st):

ok... if you know x to find z subsitute it into either of the equations that only contain x and z... e.g. 5x -2z = 9 then when you get z... substitute x and z into any of the original equations to find y

OpenStudy (ammarah):

i got 1,1.-2

OpenStudy (campbell_st):

I think you need to go and recheck y.

OpenStudy (wolf1728):

One of those ('y') is incorrect.

OpenStudy (ammarah):

y=6?

OpenStudy (campbell_st):

thats correct,

OpenStudy (ammarah):

ok i have another one please x-y+z=5 2y+3z=14 =3y+2z=5

OpenStudy (campbell_st):

sure... what do you suggest.. as the 1st step

OpenStudy (campbell_st):

is the last equation -3y +2z = 5

OpenStudy (ammarah):

yes